(xz+yz)dx/dy + (xz-yz)dz/dy = x2+y2 find the general solution


1
Expert's answer
2021-01-28T04:40:44-0500

(xz+yz)p+(xz-yz)q=x2+y2

P=xz+yz

Q=xz-yz

R=x2+y2

"\frac{dx}{P}= \frac{dy}{Q}=\frac{dz}{R}"

"\frac{dx}{xz+yz}= \frac{dy}{xz-yz}=\frac{dz}{x^2+y^2}"

Multipliers: -x,y,z

-xdx+ydy+zdz=0

1/2(-x2+y2+z2)=C1

Multipliers: x,-y,-z

xdx-ydy-zdz=0

1/2(x2-y2-z2)=C2

"\phi(c_{1},c_{2})=0"

"\phi( 1/2(-x^2+y^2+z^2),1/2(x^2-y^2-z^2))=0"